OpenStudy (mathmusician):

AP Calculus Help Please What is the area bounded by y = x^2 and y =3x?

OpenStudy (mathmusician):

@freckles

OpenStudy (freckles):

so let's look at the intersections of these to curves or lines

OpenStudy (freckles):

To find when y=x^2 and when y=3x intersect, one way to go about this is to find for what x value is their y values the same... The equations are already solved for y... so we need to solve x^2=3x to find for what x values the y values are the same .

OpenStudy (mathmusician):

one intersection looks like its at (0,0) and the other looks like it is at (3,9) by looking at the graph

OpenStudy (freckles):

that is correct

OpenStudy (freckles):

so we will do a little doodle